### v1

# 3DGSR: Implicit Surface Reconstruction with 3D Gaussian Splatting

Abstract.
In this paper, we present an implicit surface reconstruction method with 3D Gaussian Splatting (3DGS), namely 3DGSR, that allows for accurate 3D reconstruction with intricate details while inheriting the high efficiency and rendering quality of 3DGS. The key insight is incorporating an implicit signed distance field (SDF) within 3D Gaussians to enable them to be aligned and jointly optimized. First, we introduce a differentiable SDF-to-opacity transformation function that converts SDF values into corresponding Gaussians’ opacities. This function connects the SDF and 3D Gaussians, allowing for unified optimization and enforcing surface constraints on the 3D Gaussians. During learning, optimizing the 3D Gaussians provides supervisory signals for SDF learning, enabling the reconstruction of intricate details. However, this only provides sparse supervisory signals to the SDF at locations occupied by Gaussians, which is insufficient for learning a continuous SDF. Then, to address this limitation, we incorporate volumetric rendering and align the rendered geometric attributes (depth, normal) with those derived from 3D Gaussians. This consistency regularization introduces supervisory signals to locations not covered by discrete 3D Gaussians, effectively eliminating redundant surfaces outside the Gaussian sampling range. Our extensive experimental results demonstrate that our 3DGSR method enables high-quality 3D surface reconstruction while preserving the efficiency and rendering quality of 3DGS. Besides, our method competes favorably with leading surface reconstruction techniques while offering a more efficient learning process and much better rendering qualities. The code will be available at https://github.com/CVMI-Lab/3DGSR.
Gaussian Splatting, Implicit Function, Signed Distance Function, Volumetric Rendering
Figure 1. Our method, called 3DGSR, achieves accurate 3D surface reconstruction with rich details, while maintaining the efficiency and high-quality rendering of 3DGS. The left part illustrates the performance comparison of 3DGSR with cutting-edge reconstruction methods. The right part shows the capability of our method to achieve high-quality reconstruction and rendering results simultaneously.
1. Introduction
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a new state-of-the-art approach for high-quality novel view synthesis. This method represents the geometry and appearance of a 3D scene as a set of Gaussians, which are then optimized from posed multi-view images. Thanks to its Gaussian rasterization pipeline, 3DGS achieves real-time efficiency, even when rendering high-resolution outputs. Despite its impressive rendering quality and efficiency, 3DGS generates only noisy, incomplete points for 3D geometry and struggles to accurately reconstruct the 3D surface of a scene (Tang et al., 2023a). These, however, are highly desired in various geometry-related applications, such as 3D reconstruction, geometry editing, animation, and relighting, among others.
This motivates us to investigate how to enable 3DGS for high-quality surface reconstruction while preserving its rendering capabilities and speed. The inability of 3DGS to faithfully represent a 3D surface stems from its unstructured point-based geometry representation (i.e., the center of Gaussian). During optimization, the positions of these points (Gaussians) are adjusted to capture the geometry of the scene. However, the learning objective is to improve the rendering quality, which may not necessarily lead to accurate positions. The limited number of views utilized for learning further exacerbates this issue. Besides, the optimization process does not consider the geometric constraints among the points that are indeed constrained by the underlying geometry (i.e., surfaces), leading to noisy points. Further, the discrete and unstructured points make it difficult to extract 3D surfaces through post-processing, such as Poisson surface reconstruction (Kazhdan and Hoppe, 2013).
In this paper, we introduce 3DGSR, a novel approach that incorporates a neural implicit signed distance field (SDF) within Gaussians for geometry modeling, which is then jointly optimized and aligned with 3D Gaussians. First, we propose a differentiable SDF-to-opacity transformation function that converts SDF values into corresponding Gaussians’ opacities, thus connecting SDF and Gaussians. This association enforces surface constraints on Gaussians, aligns the SDF with Gaussians, and facilitates their unified optimization. Next, during learning, optimizing the Gaussians will provide supervisory signals for updating the SDF, allowing the reconstruction of intricate details. However, this only offers sparse supervisory signals to the SDF at locations occupied by the Gaussians, which proves inadequate for optimizing a continuous SDF and may lead to the formation of redundant structures. To address this issue, we further integrate volumetric rendering (Mildenhall et al., 2021; Wang et al., 2021) and align the rendered depth (normal) with that obtained from 3D Gaussians through consistency loss to optimize the SDF. This approach effectively regularizes locations not covered by 3D Gaussians, eliminating unnecessary surface reconstructions beyond the Gaussian sampling range (see Fig. 3). It is worth noting that since we only sample a few rays during each learning iteration, this does not introduce significant computational costs. The coupled representation and learning of the 3D Gaussians and the implicit SDF field form a positive feedback cycle, enabling them to improve each other through mutual learning and ultimately yield high-quality rendering and surface reconstruction. As illustrated in Fig. 1, our approach achieves accurate reconstruction and high-quality rendering results simultaneously.
Comprehensive experiments conducted on various datasets show that our method is capable of producing high-quality reconstruction results while simultaneously preserving the efficiency and qualities of 3DGS. When compared to state-of-the-art methods for surface reconstruction (Wang et al., 2021; Yariv et al., 2021; Tang et al., 2023b), our approach substantially surpasses them in rendering quality and facilitates the reconstruction of intricate structures with significantly shorter learning time.
2. Related Works
Our method builds upon prior works in 3D reconstruction with multi-view images to enable fast surface reconstruction and novel view synthesis. We discuss prior works in multi-view 3D reconstruction, neural implicit representation, and point-based rendering.
Multi-view Stereo
Traditional multi-view stereo (MVS) methods have employed point-based (Barnes et al., 2009; Stereopsis, 2010; Galliani et al., 2016; Schönberger et al., 2016) and voxel-based (De Bonet and Viola, 1999; Broadhurst et al., 2001; Seitz and Dyer, 1999) representation. Point-based methods estimate the 3D positions of pixels by matching multi-view images and finding pixel correspondences. The surface can then be obtained by performing screened Poisson surface reconstruction (Kazhdan and Hoppe, 2013) on the resulting point clouds. However, these methods heavily rely on matching accuracy and struggle to reconstruct textureless regions. On the other hand, voxel-based methods cannot achieve high-quality reconstruction due to the resolution limitation caused by high memory consumption.
Neural Implicit Representations
In recent years, the fields implicitly predicted by neural networks have gained significant attention in both novel view synthesis (Lombardi et al., 2019; Sitzmann et al., 2019; Mildenhall et al., 2021) and 3D reconstruction (Yariv et al., 2020; Niemeyer et al., 2020; Kellnhofer et al., 2021; Jiang et al., 2020; Liu et al., 2020b). Neural Radiance Field (NeRF) (Mildenhall et al., 2021) has demonstrated excellent performance in high-quality novel view synthesis by modeling scenes with a density field and a view-dependent color field. However, NeRF’s density field cannot extract high-quality surfaces. To achieve accurate surface reconstruction, NeuS (Wang et al., 2021) represents the scene with a signed distance function (SDF) and derives the alpha value from the precondition that the zero-level surface contributes the most to the volume rendering. Following NeuS, Darmon et al(Darmon et al., 2022) improves the reconstruction quality with an additional warping loss that ensures photometric consistency. Neuralangelo (Li et al., 2023b) introduces multi-resolution hash grids (Müller et al., 2022) and regularizes the model with numerical gradients and coarse-to-fine training. However, these methods often require extensive training time, which can limit their practical applicability.
Several works have focused on accelerating neural implicit representation by using implicit-explicit hybrid representation with grids (Sun et al., 2022; Karnewar et al., 2022; Müller et al., 2022; Wang et al., 2023a; Fridovich-Keil et al., 2022), trees (Yu et al., 2021; Liu et al., 2020a), points (Xu et al., 2022; Dai et al., 2023; Chen et al., 2023b), or low-rank decomposition (Chan et al., 2022; Hu et al., 2023; Chen et al., 2022). These methods speed up the training process and promote the rendering speed of radiance fields to a remarkable real-time level. However, fast and high-quality surface reconstruction from multi-view images remains a challenging problem.
Point-based Rendering
Unstructured points (Gross and Pfister, 2011) can be efficiently rendered to obtain novel-view images. However, directly rasterizing colored point clouds (Grossman and Dally, 1998) with fixed point sizes can lead to disconnected artifacts such as discontinuity, holes, and aliasing. To address these issues, previous works (Botsch et al., 2005; Pfister et al., 2000; Ren et al., 2002; Zwicker et al., 2001) extend points to larger primitives (e.g., spheres, ellipsoids, and surfels) and splat them onto the image plane. Differentiable point-based methods (Rakhimov et al., 2022; Pittaluga et al., 2019; Aliev et al., 2020) tackle these artifacts by rendering a holed latent feature map and then decoding it as a full RGB image using neural networks. Alternatively, some approaches render points as primitives (e.g., surfels and Gaussians) with trainable parameters and render them differentially (Yifan et al., 2019; Zhang et al., 2022; Kerbl et al., 2023).
Most recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has demonstrated remarkable performance in terms of high-level rendering quality and real-time rendering. Subsequent works have extended 3DGS to handle 3D/4D generation (Yi et al., 2023; Tang et al., 2023a; Zielonka et al., 2023; Ren et al., 2023; Yang et al., 2023a), dynamic scenes (Luiten et al., 2023; Yang et al., 2023b, 2024; Das et al., 2023; Li et al., 2023a; Huang et al., 2023), relighting (Gao et al., 2023), and animation (Jung et al., 2023; Ye et al., 2023). Some concurrent works (Guédon and Lepetit, 2023; Chen et al., 2023a) aim to reconstruct the 3D surface from multi-view images using 3DGS. NeuSG (Chen et al., 2023a) jointly trains a NeuS (Wang et al., 2021) and 3DGS. During training, the normals derived from NeuS regularize the rotation of Gaussians, while NeuS is also regularized to ensure that the SDF values at Gaussians’ positions are close to zero. SuGaR (Guédon and Lepetit, 2023) enforces Gaussians to have limited overlap with others during the optimization of 3DGS. After that, an initial mesh can be extracted by performing Poisson reconstruction (Kazhdan and Hoppe, 2013) on Gaussians, using the shortest axis as normals. The mesh is then bound with flat Gaussians distributed on its faces and optimized to obtain refined results.
3. Preliminaries
3D Gaussian Splatting (3DGS)
3DGS (Kerbl et al., 2023) employs 3D Gaussian points, characterized by a central point $\mu$ and a covariance matrix $\Sigma$, to effectively render images from given viewpoints:
(1) $G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\right).$
To facilitate optimization, $\Sigma$ is factorized into the product of a scaling matrix $S$, represented by scale factors $s$, and a rotation matrix $R$ encoded by a quaternion $r\in\mathbb{R}^{4}$:
(2) $\Sigma=RSS^{T}R^{T},$
For rendering novel views, the 3D Gaussians are projected onto a 2D image plane according to elliptical weighted average (EWA) (Zwicker et al., 2001):
(3) $\Sigma^{\prime}=JW\Sigma W^{T}J^{T},$
where $W$ is the viewing transformation matrix, and $J$ is the Jacobian derived from projective transformations. The final color $C(u)$ at pixel $u$ is rendered by combining $N$ ordered Gaussian splats using point-based $\alpha$-blending:
(4) $C(u)=\sum_{i\in N}T_{i}\alpha_{i}c_{i},\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}).$
Here, $c_{i}$ is determined by the spherical harmonic (SH) function and the viewing direction, while $\alpha_{i}$ is computed from the 2D covariance $\Sigma^{\prime}$, modulated by the opacity $\sigma$ of each Gaussian (Kerbl et al., 2023). In addition to rendering color $c_{i}$, Eq. (4) is also employed to render normal and depth as detailed below.
Normal and Depth Estimation from Gaussian
Typically, each 3D Gaussian can be treated as a conventional point cloud by using the centers of 3D Gaussians. Then, we can leverage the estimated depth (normal) and render them into a depth (normal) map using point-based alpha blending  (Gao et al., 2023) as:
(5) $\{\mathcal{D},\mathcal{N}\}=\sum_{i\in N}T_{i}\alpha_{i}\{d_{i},\mathbf{n}_{i}\}.$
Here $d_{i}$ and $\mathbf{n}_{i}$ are the depth and normal of $i$th point respectively.
4. Method
Figure 2. Pipeline of our proposed approach for implicit surface reconstruction. We model the surface with an implicit SDF field, with which the SDF value of each 3D Gaussian can be predicted. The opacity of each Gaussian is modeled as a bell-shaped function of its distance to the implicit surface. The geometry attributes of 3D Gaussians serve as a regularization for the SDF field, while the rendered image is supervised by the captured image.
Given posed multiview images $\{\mathcal{I}_{k}\}$ and a point cloud obtained from the Structure from Motion algorithm (Schönberger and Frahm, 2016), our goal is to enable 3DGS with the ability to reconstruct accurate surface while preserving its rendering quality and efficiency. To accomplish this, we propose an approach integrating neural implicit signed distance field (SDF) within 3D Gaussians as shown in Fig. 2. A hash grid and a single-layer MLP are employed to predict the SDF value at each queried location. Additionally, we introduce a differentiable SDF-to-opacity transformation function $\Phi$ to link and align SDF and 3D Gaussians. Subsequently, given a 3D Gaussian, its opacity value is determined by querying the corresponding SDF value and transforming it using the function $\Phi$. This coupled design allows the optimization of 3D Gaussians to supply supervisory signals for updating the SDF, enabling the reconstruction of intricate details. Details are unfolded in (Sec. 4.1).
However, the above is insufficient for optimizing the continuous SDF, as it can only provide supervision at locations where Gaussians are present. This results in floating artifacts in free space due to inadequate supervision (see Fig.3). To tackle this issue, we utilize volumetric rendering to generate depth and normals by querying points along the cast ray from the neural implicit SDF. Subsequently, a depth (normal) consistency loss is employed to align the rendered depth (normal) with those derived from 3D Gaussians, effectively regularizing the locations not covered by Gaussians and limiting floaters. The details are elaborated in  4.2.
4.1. Connect 3D Gaussians with Implicit SDF
The vanilla 3DGS method optimizes millions of Gaussians to accurately fit the training views, enabling photorealistic synthesis of novel views. However, its main emphasis is on fitting the target views, which does not guarantee high-quality 3D geometries. Moreover, the 3D Gaussians should be constrained by the underlying geometry (i.e., surfaces) of the 3D scene, which is absent in current methods. Consequently, 3DGS faces limitations in accurately reconstructing 3D surfaces (see Fig. 5).
Here, we introduce neural implicit SDF, denoted as $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{1}$, which maps each 3D position $x\in\mathbb{R}^{3}$ to its signed distance from the target surface. This is implemented using multi-resolution hash grids (Müller et al., 2022) in conjunction with a single-layer MLP. The surface $\mathrm{S}$ of a scene can then be derived by using the zero-level set of its SDF, defined by:
(6) $\mathrm{S}=\{x\in\mathbb{R}^{3}\>|\>f(\mathbf{x})=0\}.$
Next, to link the SDF with 3D Gaussians, we introduce a differentiable SDF-to-opacity transformation function. This function should adhere to the constraint that points closer to the surface contribute more to rendering a pixel and, therefore, have a higher opacity value. Consequently, we select a bell-shaped function $\Phi_{\beta}(f(x))$ controlled by a hyperparameter $\gamma$, expressed as:
(7) $\Phi_{\beta}(f(x))=\frac{e^{-\beta\cdot f(x)}}{(1+e^{-\beta\cdot f(x)})^{2}},$
where $\beta$ is a learnable parameter controlling the shape of the function.
This function transforms SDF values at a query location into the opacity of the corresponding Gaussian at that location. By linking the SDF with 3D Gaussians using a differentiable function, the SDF field can be updated through the optimization of 3D Gaussians by minimizing the photometric loss:
(8) $\mathcal{L}_{c}=\lambda\mathcal{L}_{\text{D-SSIM}}+(1-\lambda)\mathcal{L}_{1},$
This loss function encourages the Gaussian at a location $x$ that contributes more to rendering a pixel to have a higher opacity value $\Phi_{\beta}(f(x))$. This, in turn, encourages the SDF $f(x)$ to be closer to 00 and, therefore, nearer to the surface. The high-quality rendering capability of 3DGS allows for learning detailed image structures, which facilitates the reconstruction of detailed structures (see Fig. 5).
In addition to $\mathcal{L}_{c}$, we also impose a constraint as Eq. (9) to maintain consistency between the rendered normal $\mathcal{N}$, obtained directly from Gaussians via Eq. 5, and a pseudo normal $\hat{\mathcal{N}}$, computed from the rendered depth $\mathcal{D}$ derived from 3D Gaussians under the local planarity assumption.
(9) $\mathcal{L}_{n}=\|\mathcal{N}-\hat{\mathcal{N}}\|_{2}.$
Further, to encourage $f$ to approximate a valid signed distance function, we follow the Eikonal equation (Gropp et al., 2020) to penalize the deviation of $f$:
(10) $\mathcal{L}_{\text{eik}}=\mathbb{E}_{x}[(\|\triangledown f-1\|)^{2}].$
Additionally, we enforce that the distribution of Gaussians adheres to the surface by constraining the SDF values at their centers to be close to zero:
(11) $\mathcal{L}_{\text{pt}}=\|f(x_{g})\|_{1},$
where $x_{g}$ denotes the centers of 3D Gaussians.
4.2. Regularization with Volumetric Rendering
By integrating the SDF field with 3D Gaussians and training them jointly, as discussed in Section 4.1, we can attain high-fidelity rendering results with an optimized SDF field. However, due to the sparse and discrete nature of 3D Gaussians, the aforementioned supervision signals are only available for locations occupied by 3D Gaussians, predominantly near the surface. This is insufficient to regularize the continuous SDF, resulting in artifacts at locations not covered by Gaussians, such as floating artifacts in unoccupied free spaces, as illustrated in Fig. 3.
Figure 3. Comparison of our method with and without volumetric rendering regularization. The incorporation of volumetric rendering regularization proves to be highly effective in eliminating floating artifacts in free space where Gaussians are absent.
Figure 4. Comparison of volume rendering results and Gaussian splatting results on the TNT dataset. The resemblance demonstrates the efficacy of our learning process in aligning the rendered geometric attributes - depth and normal - derived from volumetric rendering (which is based on the learned SDF) and those obtained from 3DGS.
Since volumetric rendering employs ray-casting to render a pixel, it samples points along a ray that includes both occupied and free points. We propose incorporating volumetric rendering for optimization. The depth map and normal map of volume rendering can effectively represent the occupancy distribution in the whole space, while those of sparse Gaussians aligns well with the observed image. This inspires us to adopt the rendering results of Gaussians’ geometry attributes as the guidance of volume rendering, which can be utilized as the medium for propagating gradients. Specifically, we employ volumetric rendering to render depth (normals) and supervise them using depth (normals) derived from Gaussians using Eq. (5). This can provide regularizations to SDF values at locations uncovered by Gaussians.
More specifically, for each pixel, the ray cast from it is denoted as $\{\mathbf{p}(t)=r_{o}+t\cdot r_{d}\>\|\>t>0\}$, where $r_{o}$ is the camera center and $r_{v}$ is the unit direction vector of the ray. We accumulate the depth $\hat{\mathbf{D}}(r)$ and normal $\hat{\mathbf{N}}(r)$ along the cast ray (Wang et al., 2021):
(12) $\tilde{\mathbf{D}}(r)=\sum_{i=1}^{M}T^{r}_{i}\alpha_{i}t^{r}_{i}\;,\tilde{\mathbf{N}}(r)=\sum_{i=1}^{M}T^{r}_{i}\alpha_{i}\tilde{n}^{r}_{i}\;,$
where $\tilde{n}$ is the normal computed by the gradient of SDF field $\triangledown f$, $T^{r}_{i}$ and $\alpha_{i}$ represent the transmittance and alpha value (a.k.a opacity) of the sample point, and their values can be computed by
(13) $T^{r}_{i}=\prod_{j=1}^{i-1}(1-\alpha_{i}),~{}\alpha_{i}=\max\left(\frac{\phi_{s}(f(x_{i}))-\phi_{s}(f(x_{i+1}))}{\phi_{s}(f(x_{i}))},0\right),\;$
where $\delta^{r}_{i}$ is the distance between adjacent sample points and $\phi_{s}$ is the Sigmoid function.
After that, we utilize the depth $\mathcal{D}$ and normal $\mathcal{N}$ rendered by Gaussian Splatting to supervise the volumetric rendering results $\mathbf{D}$ and $\mathbf{N}$ by optimizing the following objectives:
(14) $\mathcal{L}_{\text{vd}}=\sum_{\mathbf{r}\in\mathcal{R}}\|\mathcal{D}(r)-\tilde{\mathbf{D}}(r)\|_{2}\;,$
and
(15) $\mathcal{L}_{\text{vn}}=\sum_{\mathbf{r}\in\mathcal{R}}\|{\mathcal{N}}(\mathbf{r})-\tilde{\mathbf{N}}(\mathbf{r})\|_{1}+\|1-\mathcal{N}(\mathbf{r})\tilde{\mathbf{N}}(\mathbf{r})\|_{1}\;,$
where $\mathcal{R}$ is the union of all training rays. These constraints regularize all the sampling points along the training rays, resulting in dense supervision for the implicit field and eliminating floating artifacts. As shown in Fig. 2, there exist noisy SDF distributions (negative value) floating over the free space between the camera and the object, which is then removed by applying the regularization with volumetric rendering. Accurate geometry can be obtained through our careful training strategy and model design. We visualize the optimized depth map and normal map of both volume rendering and Gaussian rendering in Fig 4.
The overall training objective is defined as a combination of all aforementioned losses:
(16) $\mathcal{L}=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{n}+\lambda_{2}\mathcal{L}_{\text{eik}}+\lambda_{3}\mathcal{L}_{\text{pt}}+\lambda_{4}\mathcal{L}_{\text{vd}}+\lambda_{5}\mathcal{L}_{\text{vn}},$
where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}$ are hyperparameters that control the relative importance of each loss term.
5. Experiments
5.1. Experimental Settings
Datasets
Our evaluation encompasses three datasets: the NeRF synthetic dataset (Mildenhall et al., 2021), the DTU real-captured dataset (Jensen et al., 2014), and the Tanks&Temples (TNT) large-scale dataset (Knapitsch et al., 2017). The NeRF synthetic dataset consists of eight objects, each rendered via path tracing to generate 100 views. For testing, we select an eighth of these images, allocating the remainder for training. The DTU and TNT datasets are processed for camera poses as per (Gao et al., 2023), with image and mask processing following the methodology outlined in (Liu et al., 2020a). The image resolutions for the DTU and TNT datasets are $1600\times 1200$ and $1920\times 1080$ pixels, respectively.
Comparisons
We evaluated our method from both surface reconstruction quality and rendering quality of novel-view synthesis. We benchmark our algorithm against several state-of-the-art methods, including NeRF (Mildenhall et al., 2021), NeuS (Wang et al., 2021), NeuSG (Chen et al., 2023a), RelighableGaussian (Gao et al., 2023), and NeRF2Mesh (Tang et al., 2023b), Neuralangelo (Li et al., 2023b), PET-NeuS (Wang et al., 2023b), BakedSDF (Yariv et al., 2023), VolSDF (Yariv et al., 2021). The quality of image synthesis is quantified using the Peak Signal-to-Noise Ratio (PSNR). For surface reconstruction accuracy, we employ the Chamfer distance and the F1 score as our metrics (Knapitsch et al., 2017). Furthermore, we also compare the rendering quality of our method with some works that solely focus on novel view synthesis tasks (Tri-MipRF (Hu et al., 2023), Zip-NeRF (Barron et al., 2023), 3DGS (Kerbl et al., 2023)).
Chair Drums Ficus Hotdog Lego Materials Mic Ship Avg
PSNR NeRF (Mildenhall et al., 2021) 34.17 25.08 30.39 36.82 33.31 30.03 34.78 29.30 31.74
Ins-NGP (Müller et al., 2022) 35.00 26.02 33.51 37.40 36.39 29.78 36.22 31.10 33.18
Mip-NeRF (Barron et al., 2021) 35.14 25.48 33.29 37.48 35.70 30.71 36.51 30.41 33.09
3D-GS (Kerbl et al., 2023) 35.36 26.15 34.87 37.72 35.78 30.00 35.36 30.80 33.32
NeuS (Wang et al., 2021) 31.22 24.85 27.38 36.04 34.06 29.59 31.56 26.94 30.20
NeRO (Liu et al., 2023) 28.74 24.88 28.38 32.13 25.66 24.85 28.64 26.55 27.48
BakedSDF (Yariv et al., 2023) 31.65 20.71 26.33 36.38 32.69 30.48 31.52 27.55 29.66
NeRF2Mesh (Tang et al., 2023b) 34.25 25.04 30.08 35.70 34.90 26.26 32.63 29.47 30.88
Ours 34.85 26.08 35.17 36.88 34.90 30.03 36.44 31.48 33.23
C-$\mathcal{L}_{1}$ VolSDF (Yariv et al., 2021) 1.18 4.03 3.01 3.22 2.26 1.57 1.13 6.42 2.86
NeuS (Wang et al., 2021) 3.99 1.27 0.94 2.12 2.56 1.39 1.00 5.38 2.33
NeRO (Liu et al., 2023) 1.27 1.97 1.22 1.88 1.90 1.33 0.87 4.95 1.92
BakedSDF (Yariv et al., 2023) 1.83 1.43 1.09 1.68 1.13 1.36 0.84 3.88 1.66
NeRF2Mesh (Tang et al., 2023b) 1.62 1.11 0.65 2.73 1.93 1.42 0.78 2.20 1.55
RelightableGaussian (Gao et al., 2023) 3.65 2.34 1.26 3.11 1.63 1.35 1.76 3.35 2.31
Ours 1.01 0.95 0.69 1.92 1.35 1.35 1.15 3.35 1.50
Table 1. We assess the quality of synthesized images and the accuracy of surface reconstruction, with each cell colored to indicate the best and second best. Our method is compared against various state-of-the-art (SOTA) approaches in tasks of novel view synthesis and surface reconstruction. It outperforms all competitors in both tasks, achieving the highest PSNR and the lowest Chamfer-$\mathcal{L}_{1}$ (C-$\mathcal{L}_{1}$) distance.
Scan ID 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
COLMAP 0.81 2.05 0.73 1.22 1.79 1.58 1.02 3.05 1.40 2.05 1.00 1.32 0.49 0.78 1.17 1.36
NeRF (Mildenhall et al., 2021) 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49
NeuS (Wang et al., 2021) 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84
VolSDF (Yariv et al., 2021) 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86
ReGS (Gao et al., 2023) 1.21 1.00 1.10 0.87 1.03 1.51 1.51 1.06 1.63 0.97 1.36 1.21 0.94 1.38 1.26 1.20
Ours 0.68 0.84 0.70 0.39 1.16 0.87 0.77 1.48 1.26 0.87 0.69 0.80 0.42 0.64 0.60 0.81
Table 2. Quantitative assessment on the DTU dataset with each cell colored to indicate the best and second best. Our method achieves the highest average quality of surface reconstruction and the lowest Chamfer-$\mathcal{L}_{1}$ distance with real-world data.
Figure 5. Qualitative comparisons of reconstructed results between NeRF2Mesh (Tang et al., 2023b), Relightable 3D Gaussian (Gao et al., 2023), and our method on the NeRF Synthetic dataset. Our method yields the highest-quality surface reconstruction preserving fine details and intricate structures.
Figure 6. Qualitative comparisons of reconstructed results between VolSDF (Yariv et al., 2021), Relightable 3D Gaussian (Gao et al., 2023), and our method on the DTU dataset. Our method yields the highest-quality surface reconstruction preserving fine details and intricate structures.
Implementation Details
Our work is implemented by PyTorch and CUDA and optimized by AdamW optimizer. We bind the Gaussian points to the SDF field for the whole training process to ensure that the Gaussian rendering process can provide a coarse SDF field, that can boost the converging speed of volumetric rendering. To ensure stable optimization, we introduce the warm-up strategy to make sure the dynamic Gaussian points can move around to the surface and provide an accurate supervised signal. The warm-up steps are set as 3000 for the Nerf synthetic dataset and 5000 for the TNT dataset. During the volumetric rendering stage, we sample $N=1024$ rays per step.
For the synthetic data, we initialize the implicit function as a sphere, and the position of Gaussian is determined by the zero-level surface of the SDF field. For the real-world scenes, which frequently exhibit intricate geometry and noised pose, we adopt the SFM points as the initialized Gaussian position, and the network is still initialized as a sphere. Furthermore, we also apply the sparse MVS points to ensure the consistency between the rendered depth $\mathcal{D}$ and projected depth map $D_{mvs}$ with $\mathcal{L}_{1}$ loss. Our models are trained for 30000 iterations, and all the experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU.
5.2. Ablation Study
We conduct ablation studies on the NeRF Synthetic dataset in Table 3 to evaluate the effectiveness of each module in our method. Six different configurations are investigated to train our model.
Volumetric Constraints.
The volumetric constraints have a minimal impact on novel view synthesis (PSNR: 32.83-33.32) yet significantly influence the reconstruction results (Chamfer-L1: 3.63-1.51). Both 3DGS rasterization and ray-based volumetric rendering rely on image supervision. However, the depth and normal representations derived from these dual methods exhibit distinct characteristics. While 3DGS excels at capturing high-frequency signals, it often suffers from floaters. Conversely, ray-based volumetric rendering yields smoother and more stable outcomes, but may struggle to capture fine details.
These two methods can complement each other and help address their respective weaknesses. Redundant Gaussian points can be effectively eliminated, and simultaneously, the high-frequency information learned by 3DGS can be employed to enhance the quality of the SDF field. This is demonstrated in Table 1 and Figure 3 of the main paper.
Learnable $\beta$.
By making $\beta$ trainable results, the conversion from SDF to 3DGS opacity can be learned automatically, which results in a 0.20 and 0.44 improvement in rendering and reconstruction performance, respectively.
Normal Supervision.
If we use the pseudo normal to supervise the normal of Gaussian points, the model will get 0.23 and 0.05 improvement in Chamfer-$\mathcal{L}_{1}$ and PSNR, respectively.
Eikonal Loss.
Removing the eikonal loss (Eq. (10)) will relax the constraint of SDF field, which can damage the surface learning and lead to worse reconstruction results. Note that better rendering quality (PSNR 33.4) is obtained in this configuration, which indicates the relaxation of geometric constraints in specific ways can aid in improving rendering quality.
Surface Alignment.
Constraining the Gaussian points close to the zero-level surface will inject more high-frequency information into the SDF field leading to detailed surface extracted results.
Reconstruction Strategy.
Compared with the marching cube with the SDF field, we also employ Poisson reconstruction to extract the surface from Gaussians with normals derived from the implicit SDF field, which exhibits a lower performance. This can be attributed to the noisy, sparse distribution of Gaussians that significantly impact the surface reconstruction quality. In contrast, the implicit SDF field provides a continuous representation, allowing for easy extraction of high-fidelity surfaces.
Method Chamfer-$\mathcal{L}_{1}$ F-score PSNR LPIPS SSIM
w/o Geo 3.63 80.20 32.83 0.0380 0.963
static $\beta$ 1.95 90.20 33.02 0.0385 0.966
w/o eq (9) 1.74 91.36 33.18 0.0383 0.967
w/o eq (10) 2.00 87.89 33.40 0.0350 0.968
w/o eq (11) 1.69 91.70 33.06 0.0388 0.966
Poisson 2.03 87.86 - - -
Ours (full) 1.51 93.50 33.23 0.0340 0.968
Table 3. Ablation study on the NeRF Synthetic dataset, where we evaluate the effect of each proposed component.
5.3. Experimental Results
Tables 1 and 2 present the quantitative results for synthetic and real-world datasets, respectively. On synthetic data, Table 1 demonstrates that our method surpasses other state-of-the-art methods in terms of the reconstruction quality, as measured by Chamfer-$\mathcal{L}_{1}$ distance. Furthermore, our method also excels in synthesizing high-quality novel view images, which are comparable to state-of-the-art novel view synthesis methods. This highlights our method’s superiority over surface reconstruction techniques and its competitive performance in novel view synthesis tasks. On real-captured DTU datasets, as shown in Table 2, our method also exhibits the best performance in terms of average surface reconstruction quality.
Besides, we also found that better reconstruction results have less contribution to the rendering quality compared to the vanilla 3D-GS. The slightly lower performance is because 3DGS is optimized solely for rendering images; while our geometry constraints reduce the freedom of the optimization of 3DGS for the rendering tasks, the best solution for rendering may not necessarily be for geometry, as evidenced by performance in Table 1 (Trade-off between rendering quality and surface reconstruction). But overall, our formulation achieves the best trade-off between rendering quality and reconstruction quality.
Additionally, we provide qualitative comparisons of the reconstruction results qualitatively and present them in Figure 5. The state-of-the-art NeRF-based 3D reconstruction method NeRF2Mesh (Tang et al., 2023b) fails to preserve fine details like fine Mic textures, Lego tracks, and stripes on the Hotdog disk. RelightableGaussian (Gao et al., 2023) incorporates shading equations to determine Gaussian appearance and trains 3D Gaussians with attributes of normal and BRDF. 3D surfaces can then be obtained by applying Screened Poisson Reconstruction (Kazhdan and Hoppe, 2013) on the 3D Gaussians assigned with normal. However, Gaussians are naturally discontinuous and unstructured, resulting in broken parts on the Mic and floaters on the Hotdog disk. Due to the absence of smooth surface modeling, its reconstructed results on Lego tracks, which contain intricate structures, are noisy and inaccurate. Combining the fitting power of Gaussian Splatting and the smooth surface modeling of the implicit SDF field, our method reconstructs surfaces with geometric details accurately and smoothly.
6. Conclusion
In this paper, we introduce 3DGSR, an efficient method for high-quality surface reconstruction using 3DGS. Our approach is based on two key components: 1) the integration of neural implicit SDF and its binding with 3D Gaussians through a differentiable SDF-to-opacity transformation function, and 2) the utilization of volumetric rendering and the SDF-and-Gaussian geometry consistency regularization for SDF optimization. The first component allows for the joint optimization of SDF and 3D Gaussians, with the optimization of Gaussians providing supervision signals for learning the SDF. The second component supplies additional supervision signals for refining the SDF in areas not occupied by Gaussians, using a consistency loss that aligns depth (normal) from SDF with that derived from Gaussians. Our extensive experiments showcase the effectiveness of 3DGSR in reconstructing high-quality surfaces outperforming those obtained from state-of-the-art reconstruction pipelines, without compromising the rendering capabilities and efficiency of 3D Gaussians. Our study does have certain limitations caused by the trade-off between rendering quality and surface smoothness. In order to render high-quality images, the smoothness of the surface may be compromised in some cases involving complex texture patterns.

### v2

@@
 In this paper, we present an implicit surface reconstruction method with 3D Gaussian Splatting (3DGS), namely 3DGSR, that allows for accurate 3D reconstruction with intricate details while inheriting the high efficiency and rendering quality of 3DGS.
-The key insight is incorporating an implicit signed distance field (SDF) within 3D Gaussians to enable them to be aligned and jointly optimized.
-First, we introduce a differentiable SDF-to-opacity transformation function that converts SDF values into corresponding Gaussians’ opacities.
-This function connects the SDF and 3D Gaussians, allowing for unified optimization and enforcing surface constraints on the 3D Gaussians.
-During learning, optimizing the 3D Gaussians provides supervisory signals for SDF learning, enabling the reconstruction of intricate details.
-However, this only provides sparse supervisory signals to the SDF at locations occupied by Gaussians, which is insufficient for learning a continuous SDF.
-Then, to address this limitation, we incorporate volumetric rendering and align the rendered geometric attributes (depth, normal) with those derived from 3D Gaussians.
-This consistency regularization introduces supervisory signals to locations not covered by discrete 3D Gaussians, effectively eliminating redundant surfaces outside the Gaussian sampling range.
-Our extensive experimental results demonstrate that our 3DGSR method enables high-quality 3D surface reconstruction while preserving the efficiency and rendering quality of 3DGS.
+The key insight is to incorporate an implicit signed distance field (SDF) within 3D Gaussians for surface modeling, and to enable the alignment and joint optimization of both SDF and 3D Gaussians.
+To achieve this, we design coupling strategies that align and associate the SDF with 3D Gaussians, allowing for unified optimization and enforcing surface constraints on the 3D Gaussians.
+With alignment, optimizing the 3D Gaussians provides supervisory signals for SDF learning, enabling the reconstruction of intricate details.
+However, this only offers sparse supervisory signals to the SDF at locations occupied by Gaussians, which is insufficient for learning a continuous SDF.
+Then, to address this limitation, we incorporate volumetric rendering and align the rendered geometric attributes (depth, normal) with that derived from 3DGS.
+In sum, these two designs allow SDF and 3DGS to be aligned, jointly optimized, and mutually boosted.
+Our extensive experimental results demonstrate that our 3DGSR enables high-quality 3D surface reconstruction while preserving the efficiency and rendering quality of 3DGS.
 Besides, our method competes favorably with leading surface reconstruction techniques while offering a more efficient learning process and much better rendering qualities.
-The code will be available at https://github.com/CVMI-Lab/3DGSR.
 Gaussian Splatting, Implicit Function, Signed Distance Function, Volumetric Rendering
+journal: TOGcopyright: acmlicensedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 6article: 198publicationmonth: 12doi: 10.1145/3687952ccs: Computing methodologies Mesh geometry models
 Figure 1.
-Our method, called 3DGSR, achieves accurate 3D surface reconstruction with rich details, while maintaining the efficiency and high-quality rendering of 3DGS.
-The left part illustrates the performance comparison of 3DGSR with cutting-edge reconstruction methods.
-The right part shows the capability of our method to achieve high-quality reconstruction and rendering results simultaneously.
+Our method, called 3DGSR, achieves accurate 3D surface reconstruction with rich details while maintaining the efficiency and high-quality rendering of 3DGS.
+The left part shows the capability of our method to achieve high-quality reconstruction and rendering results simultaneously.
+2DGS (Huang et al., 2024) is the state-of-the-art Gaussian-based reconstruction method.
 1. Introduction
@@
 Thanks to its Gaussian rasterization pipeline, 3DGS achieves real-time efficiency, even when rendering high-resolution outputs.
-Despite its impressive rendering quality and efficiency, 3DGS generates only noisy, incomplete points for 3D geometry and struggles to accurately reconstruct the 3D surface of a scene (Tang et al., 2023a).
+Despite its impressive rendering quality and efficiency, 3DGS generates only noisy, incomplete point clouds for 3D geometry and struggles to accurately reconstruct the 3D surface of a scene (Tang et al., 2023a).
 These, however, are highly desired in various geometry-related applications, such as 3D reconstruction, geometry editing, animation, and relighting, among others.
 This motivates us to investigate how to enable 3DGS for high-quality surface reconstruction while preserving its rendering capabilities and speed.
-The inability of 3DGS to faithfully represent a 3D surface stems from its unstructured point-based geometry representation (i.e., the center of Gaussian).
-During optimization, the positions of these points (Gaussians) are adjusted to capture the geometry of the scene.
-However, the learning objective is to improve the rendering quality, which may not necessarily lead to accurate positions.
+The challenges faced by 3DGS in accurately representing a 3D surface stem from (1) its unstructured point-based geometry representation (i.e., the center of Gaussian), making it difficult to extract 3D surfaces through post-processing, such as Poisson surface reconstruction (Kazhdan and Hoppe, 2013); and (2) its elliptical shape, which is not naturally aligned with surface representations without proper constraints.
+Furthermore, the optimization process does not consider the geometric constraints among the points that are indeed constrained by the underlying geometry (i.e., surfaces), leading to noisy points.
 The limited number of views utilized for learning further exacerbates this issue.
-Besides, the optimization process does not consider the geometric constraints among the points that are indeed constrained by the underlying geometry (i.e., surfaces), leading to noisy points.
-Further, the discrete and unstructured points make it difficult to extract 3D surfaces through post-processing, such as Poisson surface reconstruction (Kazhdan and Hoppe, 2013).
-In this paper, we introduce 3DGSR, a novel approach that incorporates a neural implicit signed distance field (SDF) within Gaussians for geometry modeling, which is then jointly optimized and aligned with 3D Gaussians.
-First, we propose a differentiable SDF-to-opacity transformation function that converts SDF values into corresponding Gaussians’ opacities, thus connecting SDF and Gaussians.
-This association enforces surface constraints on Gaussians, aligns the SDF with Gaussians, and facilitates their unified optimization.
-Next, during learning, optimizing the Gaussians will provide supervisory signals for updating the SDF, allowing the reconstruction of intricate details.
-However, this only offers sparse supervisory signals to the SDF at locations occupied by the Gaussians, which proves inadequate for optimizing a continuous SDF and may lead to the formation of redundant structures.
-To address this issue, we further integrate volumetric rendering (Mildenhall et al., 2021; Wang et al., 2021) and align the rendered depth (normal) with that obtained from 3D Gaussians through consistency loss to optimize the SDF.
-This approach effectively regularizes locations not covered by 3D Gaussians, eliminating unnecessary surface reconstructions beyond the Gaussian sampling range (see Fig.
+Figure 2. Oversmooth reconstruction results from 2DGS (Huang et al., 2024) lead to poorer performance, reflected by a higher chamfer-$\mathcal{L}_{1}$ error (1.40) compared to 3DGSR (0.72).
+To this end, some recent works have explored the regularization of 3DGS (Guédon and Lepetit, 2023) or the use of 2DGS (Huang et al., 2024) or Gaussian surfels (Dai et al., 2024) for learning.
+Subsequently, a post-processing step, employing depth fusion (Huang et al., 2024) or Poisson reconstruction (Dai et al., 2024), is required to obtain a 3D surface.
+Despite achieving promising results, these approaches tend to sacrifice detailed structures, resulting in smoother surfaces (Fig.
+2) and suffering from limitations inherent to the reconstruction method.
+Moreover, these formulations essentially decouple surface reconstruction from GS learning, which not only prevents the surface from regularizing the learning of GS but also leads to alignment issues and the accumulation of errors.
+In this paper, we introduce 3DGSR, a simple yet effective method that integrates a lightweight neural implicit signed distance field (SDF) within Gaussians for surface modeling, which is optimized and jointly aligned with 3D Gaussians.
+We start by examining the alignment of an SDF field and 3DGS, enabling unified optimization and synchronization of the surface and appearance.
+Inspired by the intuition that Gaussians near the surface should possess a high opacity value, we explore a tight coupling strategy that uses SDF to directly control Gaussian opacity with a differentiable SDF-to-opacity transformation function.
+Although this improves reconstruction quality, we find that surface reconstruction can be disrupted by high-frequency textures, particularly in low-frequency flat surface areas, leading to noisy reconstruction (Fig.
 3).
-It is worth noting that since we only sample a few rays during each learning iteration, this does not introduce significant computational costs.
-The coupled representation and learning of the 3D Gaussians and the implicit SDF field form a positive feedback cycle, enabling them to improve each other through mutual learning and ultimately yield high-quality rendering and surface reconstruction.
-As illustrated in Fig.
-1, our approach achieves accurate reconstruction and high-quality rendering results simultaneously.
-Comprehensive experiments conducted on various datasets show that our method is capable of producing high-quality reconstruction results while simultaneously preserving the efficiency and qualities of 3DGS.
-When compared to state-of-the-art methods for surface reconstruction (Wang et al., 2021; Yariv et al., 2021; Tang et al., 2023b), our approach substantially surpasses them in rendering quality and facilitates the reconstruction of intricate structures with significantly shorter learning time.
-2. Related Works
+This issue can also be observed in the concurrent work GOF (Yu et al., 2024b) (Fig.
+3), which directly leverages Gaussian opacity and volume rendering for surface extraction.
+Therefore, instead of tight coupling SDF and 3D Gaussians, we further investigate a loose coupling strategy that only utilizes SDF to constrain the distribution and orientations of Gaussians.
+This approach forces the Gaussians to be near the surface derived from SDF and aligns their orientations (the direction of the shortest axis) with the normal of the surface.
+By doing so, we facilitate the alignment of surface and 3D Gaussians while providing flexibility for the SDF to model surface geometry and the 3D Gaussians to capture appearance details, yielding superior results in both rendering and reconstruction (Fig.
+1).
+By employing the coupling strategies, we enable the optimization of Gaussians to supply supervisory signals for the SDF.
+However, these signals are only provided at locations occupied by the Gaussians, resulting in sparse supervision for the SDF.
+This sparsity proves insufficient for optimizing a continuous SDF and may lead to the creation of redundant structures.
+To address this issue, we integrate volumetric rendering techniques (Mildenhall et al., 2021; Wang et al., 2021) and align the rendered depth (or normal) with that obtained from the 3D Gaussians through a consistency loss.
+This approach effectively regularizes the regions not covered by the 3D Gaussians.
+It is worth noting that during each learning iteration, we only sample a few rays for rendering depth and normal using a single-layer MLP.
+This not only avoids introducing significant computational costs for training but also enables fast learning of SDF from the geometric cues provided by 3DGS.
+In turn, the improved SDF helps to regularize the 3D Gaussians effectively.
+Extensive experiments conducted on multiple datasets demonstrate that our method can produce high-quality reconstruction results while maintaining the efficiency and rendering qualities of 3DGS.
+When compared to neural implicit reconstruction methods (Wang et al., 2021; Yariv et al., 2021; Tang et al., 2023b), our approach performs on par or even surpasses them while being significantly faster and achieving superior novel view synthesis results.
+In comparison to GS-based methods (Huang et al., 2024; Yu et al., 2024b), our approach outperforms them in both surface reconstruction and novel view synthesis.
+Figure 3. Imperfect reconstruction from the non-uniform distributions of Gaussians. Directly defining the surface from the Gaussian points may lead the noisy reconstruction or artifacts.
+2. Related Work
 Our method builds upon prior works in 3D reconstruction with multi-view images to enable fast surface reconstruction and novel view synthesis.
-We discuss prior works in multi-view 3D reconstruction, neural implicit representation, and point-based rendering.
-Multi-view Stereo
-Traditional multi-view stereo (MVS) methods have employed point-based (Barnes et al., 2009; Stereopsis, 2010; Galliani et al., 2016; Schönberger et al., 2016) and voxel-based (De Bonet and Viola, 1999; Broadhurst et al., 2001; Seitz and Dyer, 1999) representation.
-Point-based methods estimate the 3D positions of pixels by matching multi-view images and finding pixel correspondences.
-The surface can then be obtained by performing screened Poisson surface reconstruction (Kazhdan and Hoppe, 2013) on the resulting point clouds.
-However, these methods heavily rely on matching accuracy and struggle to reconstruct textureless regions.
-On the other hand, voxel-based methods cannot achieve high-quality reconstruction due to the resolution limitation caused by high memory consumption.
+We discuss close related prior works in neural implicit representation and 3D Gaussian splatting (3DGS).
 Neural Implicit Representations
-In recent years, the fields implicitly predicted by neural networks have gained significant attention in both novel view synthesis (Lombardi et al., 2019; Sitzmann et al., 2019; Mildenhall et al., 2021) and 3D reconstruction (Yariv et al., 2020; Niemeyer et al., 2020; Kellnhofer et al., 2021; Jiang et al., 2020; Liu et al., 2020b).
+In recent years, the fields implicitly predicted by neural networks have gained significant attention in both novel view synthesis (Lombardi et al., 2019; Sitzmann et al., 2019; Mildenhall et al., 2021) and 3D reconstruction (Yariv et al., 2020; Niemeyer et al., 2020; Kellnhofer et al., 2021; Jiang et al., 2020; Liu et al., 2020b; Yu et al., 2024a).
 Neural Radiance Field (NeRF) (Mildenhall et al., 2021) has demonstrated excellent performance in high-quality novel view synthesis by modeling scenes with a density field and a view-dependent color field.
@@
 However, these methods often require extensive training time, which can limit their practical applicability.
-Several works have focused on accelerating neural implicit representation by using implicit-explicit hybrid representation with grids (Sun et al., 2022; Karnewar et al., 2022; Müller et al., 2022; Wang et al., 2023a; Fridovich-Keil et al., 2022), trees (Yu et al., 2021; Liu et al., 2020a), points (Xu et al., 2022; Dai et al., 2023; Chen et al., 2023b), or low-rank decomposition (Chan et al., 2022; Hu et al., 2023; Chen et al., 2022).
-These methods speed up the training process and promote the rendering speed of radiance fields to a remarkable real-time level.
-However, fast and high-quality surface reconstruction from multi-view images remains a challenging problem.
-Point-based Rendering
-Unstructured points (Gross and Pfister, 2011) can be efficiently rendered to obtain novel-view images.
-However, directly rasterizing colored point clouds (Grossman and Dally, 1998) with fixed point sizes can lead to disconnected artifacts such as discontinuity, holes, and aliasing.
-To address these issues, previous works (Botsch et al., 2005; Pfister et al., 2000; Ren et al., 2002; Zwicker et al., 2001) extend points to larger primitives (e.g., spheres, ellipsoids, and surfels) and splat them onto the image plane.
-Differentiable point-based methods (Rakhimov et al., 2022; Pittaluga et al., 2019; Aliev et al., 2020) tackle these artifacts by rendering a holed latent feature map and then decoding it as a full RGB image using neural networks.
-Alternatively, some approaches render points as primitives (e.g., surfels and Gaussians) with trainable parameters and render them differentially (Yifan et al., 2019; Zhang et al., 2022; Kerbl et al., 2023).
-Most recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has demonstrated remarkable performance in terms of high-level rendering quality and real-time rendering.
+3D Gaussian Splatting.
+Most recently, 3DGS (Kerbl et al., 2023) has demonstrated remarkable performance in terms of high-level rendering quality and real-time rendering.
 Subsequent works have extended 3DGS to handle 3D/4D generation (Yi et al., 2023; Tang et al., 2023a; Zielonka et al., 2023; Ren et al., 2023; Yang et al., 2023a), dynamic scenes (Luiten et al., 2023; Yang et al., 2023b, 2024; Das et al., 2023; Li et al., 2023a; Huang et al., 2023), relighting (Gao et al., 2023), and animation (Jung et al., 2023; Ye et al., 2023).
-Some concurrent works (Guédon and Lepetit, 2023; Chen et al., 2023a) aim to reconstruct the 3D surface from multi-view images using 3DGS.
-NeuSG (Chen et al., 2023a) jointly trains a NeuS (Wang et al., 2021) and 3DGS.
-During training, the normals derived from NeuS regularize the rotation of Gaussians, while NeuS is also regularized to ensure that the SDF values at Gaussians’ positions are close to zero.
+Some concurrent works (Guédon and Lepetit, 2023; Huang et al., 2024; Yu et al., 2024b) aim to reconstruct the 3D surface from multi-view images using 3DGS.
 SuGaR (Guédon and Lepetit, 2023) enforces Gaussians to have limited overlap with others during the optimization of 3DGS.
-After that, an initial mesh can be extracted by performing Poisson reconstruction (Kazhdan and Hoppe, 2013) on Gaussians, using the shortest axis as normals.
+After that, an initial mesh can be extracted by performing Poisson reconstruction (Kazhdan and Hoppe, 2013) on Gaussians.
 The mesh is then bound with flat Gaussians distributed on its faces and optimized to obtain refined results.
+2DGS (Huang et al., 2024) represents the points as 2D surfels and leverages TSDF fusion to obtain the final mesh.
+GOF (Yu et al., 2024b) leverages ray-tracing-based volume rendering of 3D Gaussians by identifying its level.
+Different from the SuGaR and 2DGS, they extract the mesh from a trained tetrahedral grid with marching tetrahedra.
+Though these methods achieve promising reconstruction results, their surface still misses fine-grained geometry details and high-quality rendering results.
+Here, we derive 3DGSR combined with explicit Gaussian rasterization and implicit volumetric rendering to get detailed surface reconstruction and high-quality rendering results.
 3. Preliminaries
+Figure 4.
+Pipeline of our proposed approach for implicit surface reconstruction.
+We model the surface with an implicit SDF field, with which the SDF value of each 3D Gaussian can be predicted.
+We propose two different coupling strategies to make the distribution of Gaussians align with the implicit SDF field.
+The geometry attributes of 3D Gaussians serve as a regularization for the SDF field, while the rendered image is supervised by the captured image.
 3D Gaussian Splatting (3DGS)
 3DGS (Kerbl et al., 2023) employs 3D Gaussian points, characterized by a central point $\mu$ and a covariance matrix $\Sigma$, to effectively render images from given viewpoints:
-(1) $G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\right).$
+(1) $G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\right),$
+where $x$ is the point position in world space.
 To facilitate optimization, $\Sigma$ is factorized into the product of a scaling matrix $S$, represented by scale factors $s$, and a rotation matrix $R$ encoded by a quaternion $r\in\mathbb{R}^{4}$:
@@
 Normal and Depth Estimation from Gaussian
-Typically, each 3D Gaussian can be treated as a conventional point cloud by using the centers of 3D Gaussians.
-Then, we can leverage the estimated depth (normal) and render them into a depth (normal) map using point-based alpha blending  (Gao et al., 2023) as:
-(5) $\{\mathcal{D},\mathcal{N}\}=\sum_{i\in N}T_{i}\alpha_{i}\{d_{i},\mathbf{n}_{i}\}.$
-Here $d_{i}$ and $\mathbf{n}_{i}$ are the depth and normal of $i$th point respectively.
+Typically, each 3D Gaussian can be treated as a point by using the centers of 3D Gaussians.
+Then, we can leverage the estimated depth and render them into a depth map using point-based alpha blending (Gao et al., 2023) as
+(5) $\mathcal{D}=\sum_{i\in N}T_{i}\alpha_{i}d_{i},$
+where $d_{i}$ is the depth of $i$th point. We then project the rendering depth into 3D space and get the pseudo normal from the local plane.
 4. Method
-Figure 2.
-Pipeline of our proposed approach for implicit surface reconstruction.
-We model the surface with an implicit SDF field, with which the SDF value of each 3D Gaussian can be predicted.
-The opacity of each Gaussian is modeled as a bell-shaped function of its distance to the implicit surface.
-The geometry attributes of 3D Gaussians serve as a regularization for the SDF field, while the rendered image is supervised by the captured image.
-Given posed multiview images $\{\mathcal{I}_{k}\}$ and a point cloud obtained from the Structure from Motion algorithm (Schönberger and Frahm, 2016), our goal is to enable 3DGS with the ability to reconstruct accurate surface while preserving its rendering quality and efficiency.
-To accomplish this, we propose an approach integrating neural implicit signed distance field (SDF) within 3D Gaussians as shown in Fig.
-2.
-A hash grid and a single-layer MLP are employed to predict the SDF value at each queried location.
-Additionally, we introduce a differentiable SDF-to-opacity transformation function $\Phi$ to link and align SDF and 3D Gaussians.
-Subsequently, given a 3D Gaussian, its opacity value is determined by querying the corresponding SDF value and transforming it using the function $\Phi$.
-This coupled design allows the optimization of 3D Gaussians to supply supervisory signals for updating the SDF, enabling the reconstruction of intricate details.
-Details are unfolded in (Sec.
-4.1).
-However, the above is insufficient for optimizing the continuous SDF, as it can only provide supervision at locations where Gaussians are present.
-This results in floating artifacts in free space due to inadequate supervision (see Fig.3).
-To tackle this issue, we utilize volumetric rendering to generate depth and normals by querying points along the cast ray from the neural implicit SDF.
-Subsequently, a depth (normal) consistency loss is employed to align the rendered depth (normal) with those derived from 3D Gaussians, effectively regularizing the locations not covered by Gaussians and limiting floaters.
-The details are elaborated in  4.2.
-4.1. Connect 3D Gaussians with Implicit SDF
+Given posed multiview images $\{\mathcal{I}_{k}\}$ and a point cloud obtained from the Structure from Motion algorithm (Schönberger and Frahm, 2016), 3DGSR integrates an implicit signed distance field (SDF) within 3D Gaussians (Fig.
+4) to reconstruct high-quality surfaces while preserving the rendering quality and efficiency of 3DGS.
+To achieve this, 3DGSR incorporates two key designs: Alignment of Surface and 3D Gaussians (Sec.
+4.1) We investiage coupling strategies to synchronize surface and 3D Gaussians, enabling their alignment and joint optimization.
+The tight coupling strategy utilizes a differentiable SDF-to-opacity transformation function to connect the SDF field and the opacity of 3D Gaussians; and the loose coupling strategy uses surface derived from SDF to constrain the distribution of Gaussians, make them near the surface and align the orientation of 3D Gaussians with surface normals.
+By aligning 3D Gaussian and the surface, we allow unified optimization of 3D Gaussians and SDF.
+However, the supervisory signal only exists at locations occupied by 3D Gaussians and is thus not sufficient for supervising the dense SDF field.
+Thus, we propose the second design: Regularization with Volumetric Rendering (Sec.
+4.2).
+We employ volumetric rendering to render depth and normal by querying points along the cast ray from the neural implicit SDF and align them with that derived from 3D GS with a consistency regularization loss.
+This provides dense supervisory signals for learning the SDF.
+The two designs work together to enforce surface constraints on 3D Gaussians, which in turn allows 3D Gaussians to assist in optimizing the SDF, resulting in high-quality surfaces with aligned 3D Gaussians.
+Compared to previous neural implicit surface reconstruction methods (Wang et al., 2021; Yariv et al., 2021), our approach utilizes 3DGS for efficient and high-quality rendering while avoiding the dependence on time-consuming color prediction branches and volumetric rendering loss for optimization.
+Instead, it can be supervised by 3DGS and directly learn the SDF from optimized geometric clues (e.g., depth and normal) derived from 3DGS, making the training process more than 20 times faster (12 hours vs.
+30 minutes).
+In comparison to Gaussian-based reconstruction methods (Huang et al., 2024), our approach unifies the optimization of 3DGS and an implicit surface, aligning them in an end-to-end manner.
+This eliminates the need for post-processing steps such as depth fusion or Poisson reconstruction and circumvents their limitations and error accumulations, leading to improved reconstruction quality and better alignment.
+4.1. Alignment of Surfaces and Gaussians
 The vanilla 3DGS method optimizes millions of Gaussians to accurately fit the training views, enabling photorealistic synthesis of novel views.
 However, its main emphasis is on fitting the target views, which does not guarantee high-quality 3D geometries.
-Moreover, the 3D Gaussians should be constrained by the underlying geometry (i.e., surfaces) of the 3D scene, which is absent in current methods.
-Consequently, 3DGS faces limitations in accurately reconstructing 3D surfaces (see Fig.
-5).
-Here, we introduce neural implicit SDF, denoted as $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{1}$, which maps each 3D position $x\in\mathbb{R}^{3}$ to its signed distance from the target surface.
+Moreover, the 3D Gaussians should be constrained by the underlying geometry (i.e., surfaces) of the 3D scene, which is absent in existing methods.
+Here, we introduce neural implicit SDF, denoted as $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{1}$ for surface modeling, which maps each 3D position $x\in\mathbb{R}^{3}$ to its signed distance from the surface.
 This is implemented using multi-resolution hash grids (Müller et al., 2022) in conjunction with a single-layer MLP.
 The surface $\mathrm{S}$ of a scene can then be derived by using the zero-level set of its SDF, defined by:
-(6) $\mathrm{S}=\{x\in\mathbb{R}^{3}\>|\>f(\mathbf{x})=0\}.$
-Next, to link the SDF with 3D Gaussians, we introduce a differentiable SDF-to-opacity transformation function.
-This function should adhere to the constraint that points closer to the surface contribute more to rendering a pixel and, therefore, have a higher opacity value.
-Consequently, we select a bell-shaped function $\Phi_{\beta}(f(x))$ controlled by a hyperparameter $\gamma$, expressed as:
+(6) $\mathrm{S}=\{x\in\mathbb{R}^{3}\>|\>f(x)=0\}.$
+Given the SDF $f$, the next question is to connect 3DGS with the implicit SDF to enable unified surface and appearance modeling and optimization.
+Here, we explore two coupling strategies: tight coupling with a differentiable SDF-to-opacity transformation function and loose coupling with distance and orientation regularization.
+Tight Coupling Strategy.
+Intuitively, points near the surface will have large opacity values and contribute more to rendering images.
+Thus, we introduce a differentiable SDF-to-opacity transformation function to tightly couple the SDF field and 3DGS.
+Consequently, we select a bell-shaped function $\Phi_{\beta}(f(x))$ controlled by a hyperparameter $\beta$, expressed as:
 (7) $\Phi_{\beta}(f(x))=\frac{e^{-\beta\cdot f(x)}}{(1+e^{-\beta\cdot f(x)})^{2}},$
@@
 This, in turn, encourages the SDF $f(x)$ to be closer to 00 and, therefore, nearer to the surface.
-The high-quality rendering capability of 3DGS allows for learning detailed image structures, which facilitates the reconstruction of detailed structures (see Fig.
-5).
-In addition to $\mathcal{L}_{c}$, we also impose a constraint as Eq.
-(9) to maintain consistency between the rendered normal $\mathcal{N}$, obtained directly from Gaussians via Eq.
-5, and a pseudo normal $\hat{\mathcal{N}}$, computed from the rendered depth $\mathcal{D}$ derived from 3D Gaussians under the local planarity assumption.
-(9) $\mathcal{L}_{n}=\|\mathcal{N}-\hat{\mathcal{N}}\|_{2}.$
-Further, to encourage $f$ to approximate a valid signed distance function, we follow the Eikonal equation (Gropp et al., 2020) to penalize the deviation of $f$:
-(10) $\mathcal{L}_{\text{eik}}=\mathbb{E}_{x}[(\|\triangledown f-1\|)^{2}].$
-Additionally, we enforce that the distribution of Gaussians adheres to the surface by constraining the SDF values at their centers to be close to zero:
-(11) $\mathcal{L}_{\text{pt}}=\|f(x_{g})\|_{1},$
-where $x_{g}$ denotes the centers of 3D Gaussians.
+Loose Coupling.
+Although the tight coupling strategy can align Gaussian points with the implicit surface and enable unified optimization for enhanced reconstructions, it may slightly compromise rendering quality and produce unsatisfactory reconstruction results with visible artifacts (See Fig.
+3).
+Upon investigation, we find that the surface and color appearances indeed exhibit different levels of detail and frequencies, which could be a contributing factor.
+For example, as shown in Fig.
+3, a chair with flat low-frequency surface areas may require numerous non-uniformly distributed Gaussian points to represent its high-frequency textures.
+In areas with dense Gaussian points, the tight coupling strategy enforces the SDF values to be closer to zero, potentially leading to noisy surface structures with unnecessary details (see Fig.
+10) if the Gaussian locations are not accurate.
+This suggests that a tight alignment might not be the optimal solution.
+Thus, we introduce a loose coupling strategy as an alternative to the tight coupling approach.
+The central concept involves using the surface derived from the SDF to regularize the distribution of 3D Gaussians.
+This strategy ensures that 3D Gaussians are positioned near the surface and their orientations are aligned with the surface normals.
+By aligning the surface normals instead of directly constraining the opacity, we can prevent interference with the supervision needed for high-frequency appearances.
+Specifically, the direction of a 3D Gaussian at location $x_{g}$ is determined by the direction of its shortest axis of the covariance matrix, represented by $n_{g}$.
+The normal at $x_{g}$ in the SDF field is derived using $\nabla{f}(x_{g})$.
+Note we normalize $\nabla{f}(x_{g})$ to a unit vector.
+We then align the direction of the 3D Gaussian and the surface normal by minimizing the direction-aligned loss, which is defined as follows:
+(9) $\mathcal{L}_{a}=||1-n_{g}\nabla{f}(x_{g})||.$
+Additionally, we constrain the 3D Gaussians to be near the surface.
+To achieve this, we query the SDF value of each Gaussian point $x_{g}$ and calculate its nearest surface location using the gradient of the SDF field as follows:
+(10) $x_{\text{nearest}}=x_{g}+f(x_{g})\cdot\nabla{f}(x_{g}),$
+where $x_{\text{nearest}}\in R$ represent the nearest 3D points of point $x_{g}$ on the surface from SDF.
+Then, we can use $\mathcal{L}_{1}$ loss to minimize their distance.
+As shown in Fig.
+3, despite its simplicity, the loose coupling strategy yields detailed and accurate surface reconstruction results where the 3D Gaussians near the surface with their orientations aligned with surface normals.
+Moreover, this approach also helps reduce the number of Gaussians by preventing the generation of Gaussian points that are far away from the surface.
+Our experiments demonstrate that we can decrease the number of Gaussian points by 20% without compromising, or even enhancing, the rendering quality.
+If not explicitly specified, we use the loose coupling strategy for alignment by default.
 4.2. Regularization with Volumetric Rendering
-By integrating the SDF field with 3D Gaussians and training them jointly, as discussed in Section 4.1, we can attain high-fidelity rendering results with an optimized SDF field.
-However, due to the sparse and discrete nature of 3D Gaussians, the aforementioned supervision signals are only available for locations occupied by 3D Gaussians, predominantly near the surface.
-This is insufficient to regularize the continuous SDF, resulting in artifacts at locations not covered by Gaussians, such as floating artifacts in unoccupied free spaces, as illustrated in Fig.
-3.
-Figure 3.
-Comparison of our method with and without volumetric rendering regularization.
-The incorporation of volumetric rendering regularization proves to be highly effective in eliminating floating artifacts in free space where Gaussians are absent.
-Figure 4.
-Comparison of volume rendering results and Gaussian splatting results on the TNT dataset.
-The resemblance demonstrates the efficacy of our learning process in aligning the rendered geometric attributes - depth and normal - derived from volumetric rendering (which is based on the learned SDF) and those obtained from 3DGS.
-Since volumetric rendering employs ray-casting to render a pixel, it samples points along a ray that includes both occupied and free points.
-We propose incorporating volumetric rendering for optimization.
-The depth map and normal map of volume rendering can effectively represent the occupancy distribution in the whole space, while those of sparse Gaussians aligns well with the observed image.
-This inspires us to adopt the rendering results of Gaussians’ geometry attributes as the guidance of volume rendering, which can be utilized as the medium for propagating gradients.
-Specifically, we employ volumetric rendering to render depth (normals) and supervise them using depth (normals) derived from Gaussians using Eq.
-(5).
-This can provide regularizations to SDF values at locations uncovered by Gaussians.
+By integrating the SDF field with 3D Gaussians and jointly training them, as discussed in Section 4.1, we can achieve high-fidelity rendering results with an optimized SDF field.
+However, due to the sparse and discrete nature of 3D Gaussians, the supervision signals mentioned earlier are only available for locations occupied by 3D Gaussians, primarily near the surface.
+This limited availability is insufficient to regularize the continuous SDF, leading to artifacts in locations not covered by 3D Gaussians, such as floating artifacts in unoccupied free spaces, as shown in Fig.
+4.
+Volumetric rendering employs ray-casting to render a pixel, sampling points along a ray that includes both occupied and free points.
+Thus, we propose incorporating volumetric rendering to optimize the dense SDF field.
+Specifically, we use volumetric rendering to render depth and normal, which can be directly derived from the SDF field, and align them with those derived from 3DGS for optimization.
+There are two main reasons for using rendered depth and normal instead of color.
+First, rendering color would require an additional color prediction MLP network, significantly increasing the learning time.
+Second, and more importantly, learning the SDF directly from geometric clues (e.g., depth and normal) can alleviate the burden of appearance modeling, accelerate the learning process, and enhance the quality.
 More specifically, for each pixel, the ray cast from it is denoted as $\{\mathbf{p}(t)=r_{o}+t\cdot r_{d}\>\|\>t>0\}$, where $r_{o}$ is the camera center and $r_{v}$ is the unit direction vector of the ray.
@@
 (12) $T^{r}_{i}=\prod_{j=1}^{i-1}(1-\alpha_{i}),~{}\alpha_{i}=\max\left(\frac{\phi_{s}(f(x_{i}))-\phi_{s}(f(x_{i+1}))}{\phi_{s}(f(x_{i}))},0\right),\;$
-where $\delta^{r}_{i}$ is the distance between adjacent sample points and $\phi_{s}$ is the Sigmoid function.
+where $\phi_{s}$ is the Sigmoid function.
 After that, we utilize the depth $\mathcal{D}$ and normal $\mathcal{N}$ rendered by Gaussian Splatting to supervise the volumetric rendering results $\mathbf{D}$ and $\mathbf{N}$ by optimizing the following objectives:
@@
 As shown in Fig.
-2, there exist noisy SDF distributions (negative value) floating over the free space between the camera and the object, which is then removed by applying the regularization with volumetric rendering.
+4, there exist noisy SDF distributions (negative value) floating over the free space between the camera and the object, which is then removed by applying the regularization with volumetric rendering.
 Accurate geometry can be obtained through our careful training strategy and model design.
-We visualize the optimized depth map and normal map of both volume rendering and Gaussian rendering in Fig 4.
-The overall training objective is defined as a combination of all aforementioned losses:
-(16) $\mathcal{L}=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{n}+\lambda_{2}\mathcal{L}_{\text{eik}}+\lambda_{3}\mathcal{L}_{\text{pt}}+\lambda_{4}\mathcal{L}_{\text{vd}}+\lambda_{5}\mathcal{L}_{\text{vn}},$
-where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}$ are hyperparameters that control the relative importance of each loss term.
+4.3. Training Objectives and Implementation Details
+To train 3DGRS, our objective is to get high-quality rendering and reconstruction results. Thus, the final loss function is defined by
+$\mathcal{L}=\mathcal{L}_{c}+\alpha_{1}\mathcal{L}_{vd}+\alpha_{2}\mathcal{L}_{vn}+\alpha_{3}\mathcal{L}_{a}+\alpha_{4}\mathcal{L}_{eik}$
+where $\mathcal{L}_{eik}$ is the Eikonal equation (Gropp et al., 2020) to penalize the deviation of the magnitude of grad $f$ from unit length:
+(15) $\mathcal{L}_{\text{eik}}=\mathbb{E}_{x}[(\|\triangledown f-1\|)^{2}].$
+In our experiments, we set $\alpha_{1}=1.0,\alpha_{2}=0.1,\alpha_{3}=0.0001,\alpha_{4}=0.1$.
+We implement the 3DGSR with CUDA and tinycudann (Müller, 2021), building upon the framework of 3DGS.
+We customize the normal and depth maps CUDA rendering kernel.
+During training, we increase the number of 3D Gaussian primitives following the adaptive control strategy in 3DGS.
+During the volumetric rendering stage, we sample $N=1024$ rays per step.
+For the synthetic data, we initialize the implicit function as a sphere, and the position of Gaussian is randomly initialized.
+For the real-world scenes, which frequently exhibit intricate geometry and noisy poses, we adopt the SFM points as the initialized Gaussian position, and the network is still initialized as a sphere.
+Furthermore, we also apply the sparse SFM points to ensure the consistency between the rendered depth $\mathcal{D}$ and projected depth map $D_{sfm}$ with $\mathcal{L}_{2}$ loss.
+Specifically, when optimizing the pose using SFM algorithms like COLMAP, we can also obtain sparse 2D key points along with their corresponding 3D point locations.
+Given an image and its camera pose, we project visible 3D points into the camera space and use the z-value as the depth $D_{sfm}$ for the corresponding 2D key points.
+We then apply an $\mathcal{L}_{2}$ loss to enforce consistency between the rendered depth $\mathcal{D}$ and the projected sparse depth $D_{sfm}$ based on the measurements.
+Our models are trained for 30000 iterations, and all the experiments are conducted on a single NVIDIA RTX 3090 GPU.
 5. Experiments
@@
 Datasets
-Our evaluation encompasses three datasets: the NeRF synthetic dataset (Mildenhall et al., 2021), the DTU real-captured dataset (Jensen et al., 2014), and the Tanks&Temples (TNT) large-scale dataset (Knapitsch et al., 2017).
-The NeRF synthetic dataset consists of eight objects, each rendered via path tracing to generate 100 views.
-For testing, we select an eighth of these images, allocating the remainder for training.
-The DTU and TNT datasets are processed for camera poses as per (Gao et al., 2023), with image and mask processing following the methodology outlined in (Liu et al., 2020a).
-The image resolutions for the DTU and TNT datasets are $1600\times 1200$ and $1920\times 1080$ pixels, respectively.
-Comparisons
-We evaluated our method from both surface reconstruction quality and rendering quality of novel-view synthesis.
-We benchmark our algorithm against several state-of-the-art methods, including NeRF (Mildenhall et al., 2021), NeuS (Wang et al., 2021), NeuSG (Chen et al., 2023a), RelighableGaussian (Gao et al., 2023), and NeRF2Mesh (Tang et al., 2023b), Neuralangelo (Li et al., 2023b), PET-NeuS (Wang et al., 2023b), BakedSDF (Yariv et al., 2023), VolSDF (Yariv et al., 2021).
-The quality of image synthesis is quantified using the Peak Signal-to-Noise Ratio (PSNR).
-For surface reconstruction accuracy, we employ the Chamfer distance and the F1 score as our metrics (Knapitsch et al., 2017).
-Furthermore, we also compare the rendering quality of our method with some works that solely focus on novel view synthesis tasks (Tri-MipRF (Hu et al., 2023), Zip-NeRF (Barron et al., 2023), 3DGS (Kerbl et al., 2023)).
+Figure 5. Qualitative comparisons of different coupling methods and the influence of geometric constraint.
 Chair Drums Ficus Hotdog Lego Materials Mic Ship Avg
@@
 NeuS (Wang et al., 2021) 31.22 24.85 27.38 36.04 34.06 29.59 31.56 26.94 30.20
+NeuS2 (Wang et al., 2023) 31.55 24.65 29.61 34.84 31.63 27.68 34.23 28.92 30.39
 NeRO (Liu et al., 2023) 28.74 24.88 28.38 32.13 25.66 24.85 28.64 26.55 27.48
@@
 NeRF2Mesh (Tang et al., 2023b) 34.25 25.04 30.08 35.70 34.90 26.26 32.63 29.47 30.88
-Ours 34.85 26.08 35.17 36.88 34.90 30.03 36.44 31.48 33.23
+2DGS (Huang et al., 2024) 35.05 26.05 35.57 37.36 35.10 29.74 35.09 30.60 33.07
+Ours (Tight Coupling) 34.85 26.08 35.17 36.88 34.90 30.03 36.44 31.48 33.23
+Ours (Loose Coupling) 35.69 26.35 35.65 37.98 36.19 30.31 37.12 31.65 33.86
 C-$\mathcal{L}_{1}$ VolSDF (Yariv et al., 2021) 1.18 4.03 3.01 3.22 2.26 1.57 1.13 6.42 2.86
@@
 RelightableGaussian (Gao et al., 2023) 3.65 2.34 1.26 3.11 1.63 1.35 1.76 3.35 2.31
-Ours 1.01 0.95 0.69 1.92 1.35 1.35 1.15 3.35 1.50
+Ours (Tight Coupling) 1.01 0.95 0.69 1.92 1.35 1.35 1.15 3.35 1.50
+Ours (Loose Coupling) 0.99 0.93 0.68 1.90 1.32 1.33 1.15 2.64 1.37
 Table 1.
@@
 It outperforms all competitors in both tasks, achieving the highest PSNR and the lowest Chamfer-$\mathcal{L}_{1}$ (C-$\mathcal{L}_{1}$) distance.
+Our evaluation encompasses three datasets: the NeRF synthetic dataset (Mildenhall et al., 2021), the DTU real-captured dataset (Jensen et al., 2014), and the Tanks & Temples (TNT) large-scale dataset (Knapitsch et al., 2017).
+The NeRF synthetic dataset consists of eight objects, each rendered via path tracing to generate 100 views.
+The DTU and TNT datasets are processed to obtain camera poses in accordance with (Gao et al., 2023), while the images and masks are processed following the methodology outlined in (Liu et al., 2020a).
+The image resolutions for the DTU and TNT datasets are $800\times 600$ and $1920\times 1080$, respectively.
+Comparisons
+We benchmark our method on two tasks: surface reconstruction and novel view synthesis.
+Our approach is compared against several state-of-the-art 1) Neural surface reconstruction methods, including NeuS (Wang et al., 2021), NeuSG (Chen et al., 2023), RelightableGaussian (Gao et al., 2023), NeRF2Mesh (Tang et al., 2023b), Neuralangelo (Li et al., 2023b), NeUS2 (Wang et al., 2023), BakedSDF (Yariv et al., 2023), VolSDF (Yariv et al., 2021), and 2) Gaussian-based reconstruction methods such as SuGaR (Guédon and Lepetit, 2023), 2DGS (Huang et al., 2024), and GOF (Yu et al., 2024b).
+It is worth noting that GOF (Yu et al., 2024b) is a concurrent work.
+Additionally, we also compare the novel view synthesis capabilities of our method with works that focus solely on novel view synthesis tasks, including NeRF (Mildenhall et al., 2021), Mip-NeRF (Barron et al., 2021), Ins-NGP (Müller et al., 2022), NeUS2 (Wang et al., 2023), and 3DGS (Kerbl et al., 2023).
+The quality of image synthesis is quantified using PSNR.
+For surface reconstruction accuracy, we employ Chamfer distance and the F1 score (Knapitsch et al., 2017).
 Scan ID 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
-COLMAP 0.81 2.05 0.73 1.22 1.79 1.58 1.02 3.05 1.40 2.05 1.00 1.32 0.49 0.78 1.17 1.36
 NeRF (Mildenhall et al., 2021) 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49
@@
 VolSDF (Yariv et al., 2021) 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86
+NeuS2 (Wang et al., 2023) 0.56 0.76 0.49 0.37 0.92 0.71 0.76 1.22 1.08 0.63 0.59 0.89 0.40 0.48 0.55 0.70
+Neuralangelo (Li et al., 2023b) 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43 0.61
 ReGS (Gao et al., 2023) 1.21 1.00 1.10 0.87 1.03 1.51 1.51 1.06 1.63 0.97 1.36 1.21 0.94 1.38 1.26 1.20
-Ours 0.68 0.84 0.70 0.39 1.16 0.87 0.77 1.48 1.26 0.87 0.69 0.80 0.42 0.64 0.60 0.81
+SuGaR (Guédon and Lepetit, 2023) 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
+2DGS (Huang et al., 2024) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
+GOF (Yu et al., 2024b) (concurrent) 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
+Ours 0.44 0.96 0.40 0.36 1.02 0.80 0.64 1.20 1.08 0.97 0.54 0.72 0.37 0.52 0.42 0.70
 Table 2.
-Quantitative assessment on the DTU dataset with each cell colored to indicate the best and second best.
+Quantitative assessment on the DTU dataset with each cell colored to indicate the best second best and third best.
 Our method achieves the highest average quality of surface reconstruction and the lowest Chamfer-$\mathcal{L}_{1}$ distance with real-world data.
 Figure 6.
-Qualitative comparisons of reconstructed results between NeRF2Mesh (Tang et al., 2023b), Relightable 3D Gaussian (Gao et al., 2023), and our method on the NeRF Synthetic dataset.
+Qualitative comparisons of reconstructed results between Relightable 3D Gaussian (Gao et al., 2023), 2DGS (Huang et al., 2024) and our method on the DTU dataset.
 Our method yields the highest-quality surface reconstruction preserving fine details and intricate structures.
-Figure 6.
-Qualitative comparisons of reconstructed results between VolSDF (Yariv et al., 2021), Relightable 3D Gaussian (Gao et al., 2023), and our method on the DTU dataset.
-Our method yields the highest-quality surface reconstruction preserving fine details and intricate structures.
-Implementation Details
-Our work is implemented by PyTorch and CUDA and optimized by AdamW optimizer.
-We bind the Gaussian points to the SDF field for the whole training process to ensure that the Gaussian rendering process can provide a coarse SDF field, that can boost the converging speed of volumetric rendering.
-To ensure stable optimization, we introduce the warm-up strategy to make sure the dynamic Gaussian points can move around to the surface and provide an accurate supervised signal.
-The warm-up steps are set as 3000 for the Nerf synthetic dataset and 5000 for the TNT dataset.
-During the volumetric rendering stage, we sample $N=1024$ rays per step.
-For the synthetic data, we initialize the implicit function as a sphere, and the position of Gaussian is determined by the zero-level surface of the SDF field.
-For the real-world scenes, which frequently exhibit intricate geometry and noised pose, we adopt the SFM points as the initialized Gaussian position, and the network is still initialized as a sphere.
-Furthermore, we also apply the sparse MVS points to ensure the consistency between the rendered depth $\mathcal{D}$ and projected depth map $D_{mvs}$ with $\mathcal{L}_{1}$ loss.
-Our models are trained for 30000 iterations, and all the experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU.
-5.2. Ablation Study
-We conduct ablation studies on the NeRF Synthetic dataset in Table 3 to evaluate the effectiveness of each module in our method. Six different configurations are investigated to train our model.
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+NeRF (Barron et al., 2023) 31.74 0.947 0.056
+NeUS2 (Wang et al., 2023) 30.39 0.944 0.064
+3D-GS (Kerbl et al., 2023) 33.32 0.970 0.031
+2DGS (Huang et al., 2024) 33.07 0.967 0.033
+Ours 33.86 0.970 0.031
+Table 3. More quantitative assessment on the NeRF synthetic dataset.
+5.2. Main Results and Comparisons
+Quantitative analysis
+Tables 12 and 3 present the quantitative results for NeRF-synthetic and DTU datasets, respectively.
+On the NeRF-synthetic dataset (Table 1 and Table 3), our approach achieves the best performance in both 3D surface reconstruction and novel view synthesis.
+Notably, our approach even outperforms the leading neural surface reconstruction methods for reconstruction evaluation while being much faster in training and yielding much better novel view synthesis results.
+Besides, our approach also delivers higher novel-view synthesis results than 3DGS (Kerbl et al., 2023) and 2DGS (Huang et al., 2024) in terms of PSNR, while achieving performance comparable to 3DGS on SSIM and LPIPS.
+On the DTU dataset (Table 2), our method also exhibits the best performance in terms of average surface reconstruction quality when compared with Gaussian-based reconstruction methods.
+It also performs favorably when compared with leading neural implicit reconstruction methods, which take much longer time to train and often yield inferior rendering qualities.
+We find that existing methods that achieve better reconstruction qualities tend to sacrifice the rendering qualities with lower performance than the NeRF baseline.
+This might be attributed to the different levels of detail required by surface reconstruction and novel view synthesis.
+Overall, our 3DGSR achieves excellent performance in both surface reconstruction and novel view synthesis.
+Qualitative analysis
+We provide qualitative comparisons of the reconstruction results in Figures 6 and 8.
+The state-of-the-art NeRF-based 3D reconstruction method, NeuS (Wang et al., 2021), fails to preserve fine details such as intricate Mic textures, Lego tracks, and stripes on the Hotdog disk.
+The 2DGS method (Huang et al., 2024), which utilizes 2D Gaussians and relies on an additional post-processing depth fusion step to obtain 3D surfaces, tends to produce over-smoothed reconstruction results, as evidenced by the smoothed Lego track and house.
+GOF (Yu et al., 2024b) employs Gaussian opacities and volumetric rendering to extract meshes directly from 3DGS, unifying geometry and appearance modeling for 3DGS.
+However, due to the inherent differences in the level of detail for geometry and appearance modeling, their results tend to produce noisy and non-existent structures needed for modeling high-frequency appearances, as seen in the noisy balls and Lego tracks in Figure 8.
+In contrast, our approach is capable of producing high-quality reconstructions, effectively preserving the finer details and maintaining the overall structure of the objects.
+More qualitative results and comparisons are included in Figure 9 and 10.
+Speed comparison
+We compared the training and inference speed of our method with Neuralangelo (Li et al., 2023b) and 2DGS (Huang et al., 2024), as shown in Table 4.
+When training on the DTU dataset, neural surface reconstruction methods like Neuralangelo (Li et al., 2023b) require over 30 hours to produce the final results, while Gaussian-splatting-based methods like 2DGS (Huang et al., 2024) take only 18.8 minutes.
+Our method, 3DGSR, takes 30 minutes-slightly longer than 2DGS (Huang et al., 2024), but significantly faster than Neuralangelo (Li et al., 2023b).
+This discrepancy arises because 3DGSR incorporates geometric constraints that effectively optimize the implicit representation, though querying the rays during training incurs additional time costs.
+However, when rendering an 800x600 image on the DTU dataset, both 2DGS (Huang et al., 2024) and our method achieve a frame rate of 175 fps.
+This is because surface regularization is applied solely during the training phase, allowing for the simultaneous learning of high-quality surfaces and Gaussians.
+Once the optimization is complete, the surface can be extracted offline and reused for future applications.
+Method Training Speed Inference
+Neuralangelo (Li et al., 2023b) ¿ 30 h ¡ 1 FPS
+2DGS (Huang et al., 2024) 18.8 min 175 FPS
+Ours 30 min 175 FPS
+Table 4. Training and inference speed comparison on DTU dataset.
+Method C-$\mathcal{L}_{1}$ PSNR LPIPS SSIM
+static $\beta$ (Tight coupling) 1.95 33.02 0.0385 0.966
+Poisson (Tight coupling) 2.03 - - -
+Ours (Tight coupling) 1.50 33.23 0.0340 0.968
+Ours (Loose coupling) 1.37 33.86 0.0330 0.971
+w/o depth,normal 3.63 32.83 0.0380 0.963
+w/o normal 1.74 33.18 0.0383 0.967
+Table 5.
+Ablation study of the surface and GS alignment on the NeRF Synthetic dataset and the volumetric rendering depth (normal) loss, where we evaluate the effect of each proposed component.
+C-$\mathcal{L}_{1}$ is an abbreviation for the Chamfer-$\mathcal{L}_{1}$ distance.
+5.3. Ablation Studies
+We conduct ablation studies on the NeRF Synthetic dataset in Table 5 to evaluate the effectiveness of each module in our method.
+Four different configurations are investigated to train our model.
+We also show the qualitative results in Fig.
+10.
 Volumetric Constraints.
-The volumetric constraints have a minimal impact on novel view synthesis (PSNR: 32.83-33.32) yet significantly influence the reconstruction results (Chamfer-L1: 3.63-1.51).
-Both 3DGS rasterization and ray-based volumetric rendering rely on image supervision.
-However, the depth and normal representations derived from these dual methods exhibit distinct characteristics.
-While 3DGS excels at capturing high-frequency signals, it often suffers from floaters.
-Conversely, ray-based volumetric rendering yields smoother and more stable outcomes, but may struggle to capture fine details.
-These two methods can complement each other and help address their respective weaknesses.
-Redundant Gaussian points can be effectively eliminated, and simultaneously, the high-frequency information learned by 3DGS can be employed to enhance the quality of the SDF field.
-This is demonstrated in Table 1 and Figure 3 of the main paper.
-Learnable $\beta$.
-By making $\beta$ trainable results, the conversion from SDF to 3DGS opacity can be learned automatically, which results in a 0.20 and 0.44 improvement in rendering and reconstruction performance, respectively.
-Normal Supervision.
-If we use the pseudo normal to supervise the normal of Gaussian points, the model will get 0.23 and 0.05 improvement in Chamfer-$\mathcal{L}_{1}$ and PSNR, respectively.
-Eikonal Loss.
-Removing the eikonal loss (Eq.
-(10)) will relax the constraint of SDF field, which can damage the surface learning and lead to worse reconstruction results.
-Note that better rendering quality (PSNR 33.4) is obtained in this configuration, which indicates the relaxation of geometric constraints in specific ways can aid in improving rendering quality.
-Surface Alignment.
-Constraining the Gaussian points close to the zero-level surface will inject more high-frequency information into the SDF field leading to detailed surface extracted results.
+The volumetric constraints have a minimal impact on novel view synthesis (PSNR: 32.83-33.86) but significantly influence the reconstruction results (Chamfer-$\mathcal{L}_{1}$: 3.63-1.37).
+This is because the volumetric rendering for depth and normal supervision can provide dense supervisory signals for learning the continuous SDF.
+Without these constraints, numerous noisy structures would be generated in areas not occupied by Gaussians, negatively affecting the overall reconstruction quality.
+As shown in Fig.
+10, if we directly use the tight coupling strategy without any geometry constraint, we can only get the coarse result.
 Reconstruction Strategy.
-Compared with the marching cube with the SDF field, we also employ Poisson reconstruction to extract the surface from Gaussians with normals derived from the implicit SDF field, which exhibits a lower performance.
-This can be attributed to the noisy, sparse distribution of Gaussians that significantly impact the surface reconstruction quality.
-In contrast, the implicit SDF field provides a continuous representation, allowing for easy extraction of high-fidelity surfaces.
-Method Chamfer-$\mathcal{L}_{1}$ F-score PSNR LPIPS SSIM
-w/o Geo 3.63 80.20 32.83 0.0380 0.963
-static $\beta$ 1.95 90.20 33.02 0.0385 0.966
-w/o eq (9) 1.74 91.36 33.18 0.0383 0.967
-w/o eq (10) 2.00 87.89 33.40 0.0350 0.968
-w/o eq (11) 1.69 91.70 33.06 0.0388 0.966
-Poisson 2.03 87.86 - - -
-Ours (full) 1.51 93.50 33.23 0.0340 0.968
-Table 3. Ablation study on the NeRF Synthetic dataset, where we evaluate the effect of each proposed component.
-5.3. Experimental Results
-Tables 1 and 2 present the quantitative results for synthetic and real-world datasets, respectively.
-On synthetic data, Table 1 demonstrates that our method surpasses other state-of-the-art methods in terms of the reconstruction quality, as measured by Chamfer-$\mathcal{L}_{1}$ distance.
-Furthermore, our method also excels in synthesizing high-quality novel view images, which are comparable to state-of-the-art novel view synthesis methods.
-This highlights our method’s superiority over surface reconstruction techniques and its competitive performance in novel view synthesis tasks.
-On real-captured DTU datasets, as shown in Table 2, our method also exhibits the best performance in terms of average surface reconstruction quality.
-Besides, we also found that better reconstruction results have less contribution to the rendering quality compared to the vanilla 3D-GS.
-The slightly lower performance is because 3DGS is optimized solely for rendering images; while our geometry constraints reduce the freedom of the optimization of 3DGS for the rendering tasks, the best solution for rendering may not necessarily be for geometry, as evidenced by performance in Table 1 (Trade-off between rendering quality and surface reconstruction).
-But overall, our formulation achieves the best trade-off between rendering quality and reconstruction quality.
-Additionally, we provide qualitative comparisons of the reconstruction results qualitatively and present them in Figure 5.
-The state-of-the-art NeRF-based 3D reconstruction method NeRF2Mesh (Tang et al., 2023b) fails to preserve fine details like fine Mic textures, Lego tracks, and stripes on the Hotdog disk.
-RelightableGaussian (Gao et al., 2023) incorporates shading equations to determine Gaussian appearance and trains 3D Gaussians with attributes of normal and BRDF.
-3D surfaces can then be obtained by applying Screened Poisson Reconstruction (Kazhdan and Hoppe, 2013) on the 3D Gaussians assigned with normal.
-However, Gaussians are naturally discontinuous and unstructured, resulting in broken parts on the Mic and floaters on the Hotdog disk.
-Due to the absence of smooth surface modeling, its reconstructed results on Lego tracks, which contain intricate structures, are noisy and inaccurate.
-Combining the fitting power of Gaussian Splatting and the smooth surface modeling of the implicit SDF field, our method reconstructs surfaces with geometric details accurately and smoothly.
+Compared with the marching cubes on the SDF field for surface extraction, we also employ Poisson reconstruction to extract surface from Gaussians with normals derived from the implicit SDF, which yields a much lower performance.
+This demonstrates the advantage of using learned implicit SDF to represent the surface that is continuous, more tolerant to noisy supervisions, and allows for easy extraction of high-fidelity surfaces.
+Notably, the surface is modeled as a hash grid and a single-layer MLP, which is lightweight and does not introduce many burdens for computation.
+Coupling Strategy.
+We experiment with the tight coupling and loose coupling strategies proposed by our approach.
+For the tight coupling strategy, the learning of the parameter $\beta$ impacts our performance.
+In contrast, the loose coupling strategy leads to improved results in both reconstruction and novel view synthesis, demonstrating its effectiveness in balancing the different aspects of the modeling process.
+Figure 7. Sparse point cloud and the influence of projected depth supervision $D_{sfm}$ in 2DGS and 3DGSR.
+Projected Sparse Depth $D_{sfm}$.
+Method $D_{sfm}$ Chamfer-$\mathcal{L}_{1}$
+2DGS (Huang et al., 2024) Yes 0.85
+2DGS (Huang et al., 2024) No 0.80
+Ours Yes 0.70
+Ours No 0.81
+Table 6. Ablation study of sparse depth supervision. Yes indicates the use of $D_{sfm}$ as a supervision signal, while No indicates that $D_{sfm}$ is not used for supervision.
+To ensure a fair comparison, we conducted experiments to analyze the influence of the supervision by projected sparse depth $D_{sfm}$ on both 3DGSR and 2DGS (Huang et al., 2024), presenting the quantitative results in Table 6 and qualitative results in Fig.
+7.
+As shown in Table 6, we observed that adding sparse depth cues negatively affects the performance of 2DGS (Huang et al., 2024), causing a worsening in Chamfer-$\mathcal{L}_{1}$ from 0.80 to 0.85.
+This performance drop may stem from the fact that 2DGS (Huang et al., 2024) relies on depth fusion for reconstruction, as opposed to our approach, which jointly optimizes and aligns the surface with the 3D Gaussians.
+In contrast, the addition of sparse depth supervision proves beneficial in our method, resulting in improved quantitative results (Chamfer-$\mathcal{L}_{1}$ from 0.81 to 0.70).
+These cues assist SDF optimization in regions where image supervision is sparse and camera poses are inaccurate, though they are less impactful in areas where sufficient observations are available.
+When trained using only image signals, our method can be comparable with 2DGS (Huang et al., 2024).
+Fig.
+7 further demonstrates that 2DGS (Huang et al., 2024) shows only minor improvements with sparse depth supervision, whereas our method achieves more detailed reconstructions even in the absence of such supervision.
+In conclusion, these depth cues are valuable for improving performance in cases of inaccurate camera poses and ill-posed sparse view data, especially in real-world scenes.
+Our approach can effectively utilize these free cues to enhance reconstruction quality.
 6. Conclusion
 In this paper, we introduce 3DGSR, an efficient method for high-quality surface reconstruction using 3DGS.
-Our approach is based on two key components: 1) the integration of neural implicit SDF and its binding with 3D Gaussians through a differentiable SDF-to-opacity transformation function, and 2) the utilization of volumetric rendering and the SDF-and-Gaussian geometry consistency regularization for SDF optimization.
+Our approach is based on two key components: 1) the integration of neural implicit SDF and its alignment with 3DGS, and 2) the utilization of volumetric rendering and the SDF-and-Gaussian geometry consistency regularization for SDF optimization.
 The first component allows for the joint optimization of SDF and 3D Gaussians, with the optimization of Gaussians providing supervision signals for learning the SDF.
@@
 Our extensive experiments showcase the effectiveness of 3DGSR in reconstructing high-quality surfaces outperforming those obtained from state-of-the-art reconstruction pipelines, without compromising the rendering capabilities and efficiency of 3D Gaussians.
-Our study does have certain limitations caused by the trade-off between rendering quality and surface smoothness.
-In order to render high-quality images, the smoothness of the surface may be compromised in some cases involving complex texture patterns.
+We hope our approach could inspire more future work.
+Despite its promising results, our approach still has limited capability to handle unbounded scenes, due to the use of a hash grid and implicit SDF, and also has worse results on real datasets with noisy pose estimation results.
+Furthermore, it may encounter difficulties in reconstructing transparent objects.
+These limitations highlight potential areas for further improvement and development in our method.
+7. Acknowledgement
+This work has been supported by the Hong Kong Research Grant Council - Early Career Scheme (Grant No.
+27209621), General Research Fund Scheme (Grant No.
+17202422, 17212923), Theme-based Research (Grant No.
+T45-701/22-R) and RGC Matching Fund Scheme (RMGS).
+Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.